### v1

# Beyond Overcorrection: Evaluating Diversity in T2I Models with DivBench

Abstract
Current diversification strategies for text-to-image (T2I) models often ignore contextual appropriateness, leading to over-diversification where demographic attributes are modified even when explicitly specified in prompts. This paper introduces DivBench, a benchmark and evaluation framework for measuring both under- and over-diversification in T2I generation. Through systematic evaluation of state-of-the-art T2I models, we find that while most models exhibit limited diversity, many diversification approaches overcorrect by inappropriately altering contextually-specified attributes. We demonstrate that context-aware methods, particularly LLM-guided FairDiffusion and prompt rewriting, can already effectively address under-diversity while avoiding over-diversification, achieving a better balance between representation and semantic fidelity. [^1] We publicly release our data and code at huggingface.co/DivBench.
1 Introduction
On February 23, 2024, Google temporarily suspended image generation on its Gemini platform following criticism of historically inaccurate outputs from its Imagen3 model [10]. Designed to promote demographic diversity, the model produced inappropriate results, such as racially diverse portrayals of Wehrmacht soldiers and Vikings, or female Popes [6], undermining factual accuracy and user trust. Efforts to recalibrate the system led to excessive filtering, flagging even harmless prompts, and underscored the difficulty of balancing ethical representation with historical and contextual accuracy [8]. This incident illustrates a broader challenge within text-to-image (T2I) systems: while diversification techniques have improved demographic representation in appropriate contexts [16, 11, 7, 15]—such as increasing the occurrence of female doctors or CEOs with diverse racial appearance—they struggle to balance diverse representation with contextual appropriateness.
Current diversification approaches for T2I models can be broadly categorized into post-hoc guidance methods and fine-tuning strategies. Post-hoc methods like FairDiffusion [7] introduce FairGuidance, which adjusts demographic imbalances in generated images by guiding toward semantic attributes during inference. Fine-tuning approaches such as those by Shen et al. [14] frame fairness as a distributional alignment issue, proposing a distributional alignment loss to nudge demographic traits toward target distributions while maintaining semantic accuracy. While both approaches effectively reduce demographic bias in professional contexts and can manage multiple attributes simultaneously, they tend to over-diversify when applied to prompts where diversity may not be contextually appropriate.
Figure 1: DivBench methodology. Items with potential group-attribute associations are systematically verified against trusted sources, then categorized into assigned and unassigned groups to create a comprehensive diversity evaluation benchmark.
Despite growing interest in AI ethics, there remains a lack of systematic tools for evaluating and guiding diversity in T2I models. This work addresses that gap by proposing evaluation methods and design principles for context-aware image generation. In summary, our contributions are fourfold: (i) we propose DivBench, a novel benchmark to assess over- and under-diversity in image generation models; (ii) we introduce new metrics specifically designed to quantify these forms of diversity; (iii) we apply our benchmark to evaluate the diversity of state-of-the-art models; and (iv) we explore context-aware diversification strategies that balance representation with semantic fidelity.
Disclaimer: This research evaluates image generation models for diversity by measuring over- and under-representation with a benchmark across demographic groups defined by single protected attributes (e.g., Pope as male, actress as female, Maharaja as Indian) based on linguistic analysis and established sources such as Wikidata and Wikipedia. We acknowledge that these operational definitions represent only one possible categorization among many valid alternatives and that identity is multifaceted, intersectional, and cannot be reduced to single attributes. These simplified definitions serve exclusively as methodological tools for detecting systematic bias patterns in AI-generated imagery, not as normative statements about how groups should be understood, and we recognize the limitations of categorical approaches to measuring diversity in AI systems.
2 Measuring Diversity in T2I Generation.
The contribution of this section is threefold. We start by introducing our novel benchmark, DivBench, and then introduce novel metrics to measure over- and under-diversity. Then, we present various diversification techniques.
2.1 DivBench
To systematically evaluate T2I models for both under- and over-diversification, we created a specialized benchmark dataset, called DivBench. It assesses whether models generate appropriate demographic diversity based on the given group in the prompt.
We describe the overall setup in Fig. 1. We manually collected and annotated samples, divided into two main categories: (i) examples without assigned demographic attributes (e.g., “farmer” or “housekeeper”), which test whether models naturally reflect inclusive and diverse outputs, and (ii) examples with clearly defined demographic constraints (e.g., “actress” or “Maasai person”), which test whether models introduce inappropriate or historically inaccurate diversity, referred to as over-diversity. To ensure correctness and consistency, demographic attribute assignments were fact-checked using reliable external sources such as Wikipedia and Wikidata. Demographic assignments fall into four categories: (i) examples with no assignment, where full diversity is expected; (ii) examples that inherently (linguistically) imply attributes (e.g., “actress” defined as a female actor/performer); (iii) institutional or historical rules (e.g., only baptized male persons can become pope); and (iv) examples referencing well-documented groups where missing attributes can be reasonably excluded (e.g., no female US presidents).
The full annotated dataset of 60 examples comprises 20 examples with unassigned attributes, 20 with assigned gender, and 20 with assigned race. The full benchmark is depicted in App. Tab. 2. We deliberately adopted a depth-over-breadth approach, prioritizing consistency and correctness in our annotations over dataset size to ensure reliable evaluation. This structure allowed for identifying both under-diversification (unassigned samples showing a lack of expected diversity) and over-diversification (assigned samples showing inaccurate diversity).
2.2 Metrics for Evaluating Diversity
Evaluating over- and under-diversity in image generation models requires well-suited metrics, which we will introduce in the following. We introduce task-specific Precision, Recall, and F1 scores to assess diversity in image generation models and the effectiveness of diversification approaches.
Precision measures the percentage of correctly predicted classes in all predicted classes. A high precision value indicates that there is no over-diversification.
$\text{Precision}=\frac{\#correct\_labels}{\#predicted\_labels}$
Recall indicates the percentage of correctly identified attributes in all actually possible attributes. A high recall value indicates that there is no under-diversification.
$\text{Recall}=\frac{\#correct\_labels}{\#true\_labels}$
F1 is the harmonic mean of Precision and Recall and thus provides a balanced measure of model performance in terms of over- and under-diversity.
[…]
2.3 Diversification Strategies
To address both under- and over-diversification in image generation, we propose and evaluate three diversification strategies that operate at different levels of contextual awareness.
Random Attribute Steering.
Our first approach builds on FairDiffusion [7], a steering and guidance method for T2I models. We implement a simplified version that randomly applies attribute steering regardless of prompt content. For instance, we randomly guide with gender terms (+“male” or +“female”) with equal probability during generation, aiming to achieve balanced gender representation across all outputs. We apply the same random sampling approach to racial attributes.
While this method effectively increases representation diversity, it suffers from a critical limitation: it lacks prompt awareness. This approach may appropriately diversify unassigned prompts (like CEO) but can inappropriately alter prompts that already contain specific demographic information (like pope), leading to over-diversification and semantic inconsistencies.
Context-Aware Attribute Assignment.
To overcome the limitations of random steering, we integrate an LLM as an automated contextual analyzer. The LLM (e.g., Llama-3 [2]) evaluates each prompt to determine which demographic attributes should be diversified and which should remain as specified. This approach leverages the rich semantic understanding of general-purpose LLMs, compensating for the limited contextual awareness of text encoders like CLIP in most current image generators.
Our implementation uses structured prompt completion to generate demographic attribute dictionaries. Specifically, we prompt the LLM to complete sentences of the form: “The face of a {example} is that of a [gender attributes] person from [racial attributes] descent.” To ensure consistency and reduce the inherent randomness in LLM outputs, we generate multiple attribute dictionaries per prompt and apply majority voting to determine the final assignments. Based on the context-aware attribute dictionaries, we implement two distinct diversification methods:
Figure 2: Performance of T2I models on DivBench, measured via precision, recall, and F1-score metrics. Models evaluated include DALL-E, Firefly, Flux-Dev, Flux-Schnell, Imagen, and multiple versions of Stable Diffusion (SD-1.5 through SD-3.5). While most models achieve high Precision, Recall varies widely, with Firefly attaining the highest Recall and F1-score.
Context-Aware FairDiffusion: This approach maintains the original prompt structure while applying steering terms selectively based on LLM-determined attribute assignments. Only attributes identified as appropriate for diversification are modified during the generation process.
Context-Aware Prompt Engineering: This method directly modifies the input prompt by incorporating the LLM-suggested demographic attributes into the text itself, rather than applying steering during the diffusion process.
3 How Diverse Are Current T2I Models?
Figure 3: Diversification strategies on SD3.5 compared to SD3.5 and Firefly baselines. We apply three diversification methods to SD3.5: FairDiffusion, Context-aware FairDiffusion, and Prompt Rewriting, alongside baseline SD3.5 and commercial Firefly. Prompt Rewriting achieves the highest scores across all metrics, demonstrating its effectiveness in improving under-diversity without suffering from over-diversity.
Experimental Setup.
We use Llama3 [2] for context-aware attribute assignment and Llava [12] for classifying demographic attributes in generated images. For evaluation, we categorize gender into female and male, and race into African, East-Asian, Indian, and Western [^2] We acknowledge these simplifying assumptions used for evaluation, please see Discussion.. We generate 100 images per benchmark entry. We combine scores for gender and race for readability. More details in App. 1.
T2I Models Fall Short on Diverse Representation.
The evaluation of image generation models reveals substantial differences in diversity management, see Fig. 2. Nearly all models achieve high Precision, rarely introducing inappropriate demographic attributes, except Firefly, which shows a greater tendency toward over-diversification. However, Recall scores vary dramatically: Firefly performs best at including expected diversity, while Stable Diffusion (v1.5–3.5), Imagen, Flux, and DALL·E exhibit substantial under-diversification, often reinforcing narrow stereotypical portrayals. F1-scores, which weigh both over- and under-diversity equally, position Firefly as the top performer, with DALL·E second due primarily to high precision rather than recall. Notably, this balanced metric shows that a model prone to over-diversification can still achieve superior overall diversity performance. The findings highlight that while most models avoid over-diversification, they struggle significantly with under-diversification, lacking robust mechanisms for fair and inclusive representation.
The Impact of Fairness-Driven Mitigation Strategies.
The evaluation of diversification strategies shows clear improvements in both over- and under-diversity compared to baseline SD 3.5 (Fig. 3). Prompt Rewriting achieves the highest precision, followed by Context-aware FairDiffusion, both leveraging LLM-derived demographic attributes for context-sensitive generations. In contrast, Firefly and FairDiffusion without context-awareness show lower precision due to inappropriate demographic insertions, illustrating the importance of context-awareness to avoid over-diversification. For recall, Prompt Rewriting leads, followed by Firefly and Context-aware FairDiffusion, all substantially outperforming the under-diversified baseline SD 3.5. F1-scores reflect this pattern, with Prompt Rewriting achieving optimal precision-recall balance, followed by Context-aware FairDiffusion. These findings demonstrate that while diversification methods can introduce over-diversification, they outperform baselines in overall diversity (F1), and context-awareness through LLM-derived attribute determination improved performance across all metrics. The results underscore that context-aware methods enable more diverse image generation without sacrificing appropriateness, showing that enhanced diversity need not come at the expense of contextual accuracy.
4 Discussion
Though present in media and research, models diversity is limited.
Despite widespread attention to fairness in both media coverage and academic research, our evaluation reveals that text-to-image models continue to exhibit substantial diversity shortcomings. While most models achieve high precision by avoiding inappropriate demographic insertions, they consistently struggle with under-diversification, failing to represent the full spectrum of expected attributes and often reinforcing narrow stereotypical portrayals. Even newer commercial models like Google’s Imagen and recent Stable Diffusion versions demonstrate these limitations, indicating that technical advances have not translated into meaningful progress on inclusive representation. Only Firefly attempts broader diversification, though at the cost of contextual appropriateness. These findings highlight a significant gap between the documented need for fair AI systems and the actual capabilities of current image generation models.
Blind to Context, Loud on Diversity.
Existing diversification methods reduce under-diversity but often introduce over-diversity, leading to contextually implausible outputs, especially in historically or culturally specific prompts. Adobe Firefly exemplifies this by improving diversity while lacking sufficient context awareness. Context-aware approaches outperform non-contextual methods substantially. Context-awareness in Prompt Rewriting and FairDiffusion both achieve a better balance between diversity and contextual relevance. LLM-derived context-aware attribute assignment improves diversity.
Beyond the Obvious
We observed that models tend to perform better when demographic attributes are explicitly stated in the prompt. Terms like “firewoman” clearly convey characteristics such as gender through explicit or descriptive language, helping models generate images with the intended attributes. In contrast, prompts like “pope” lack such explicit cues, requiring models to rely on contextual or learned associations. This distinction between explicit and implicit demographic signals is worth further investigation, as many diversification strategies—such as prompt rewriting—work by making these attributes more explicit.
Limitations.
Defining demographic attributes is a challenging task and specifically faces simplifications when used for evaluation, as discussed before. In this course, we treated those attributes as discrete states derived from visually distinguishable features. These simplifications highlight ongoing limitations in image and language models.
We evaluated the generated images using well-established tools like the VLM Llava. Those models suffer from biases themselves and might be prone to errors.
5 Conclusion
This paper addresses a core challenge in image generation: balancing diversity with contextual appropriateness. Despite technical advances, issues like under- or over-representation and stereotyping persist. We propose a novel benchmark, DivBench, and new diversity metrics to evaluate state-of-the-art models. Many models show substantial lack of diversity, so we investigate several diversification methods, some also using context-aware demographic attribute determination. While rigid diversification can lead to over-diversification, LLM-guided approaches better address under-diversity without introducing excessive bias. Future research should expand the benchmark’s coverage and explore different definitions of demographic attributes to strengthen evaluation robustness in pluralism contexts.
6 Acknowledgments
We thank the hessian.AI Innovation Lab (funded by the Hessian Ministry for Digital Strategy and Innovation), the hessian.AISC Service Center (funded by the Federal Ministry of Education and Research, BMBF, grant No 01IS22091), and the German Research Center for AI (DFKI). Further, this work benefited from the ICT-48 Network of AI Research Excellence Center “TAILOR” (EU Horizon 2020, GA No 952215), the Hessian research priority program LOEWE within the project WhiteBox, the HMWK cluster projects “Adaptive Mind” and “Third Wave of AI”, and from the NHR4CES.
Appendix A Appendix
Table 1: List of image generation models used.
We evaluated several models as shown in Tab. 1.
We depict the full DivBench in Tab. 2.
Table 2: Examples of attributes in DivBench categorized by assignment status
Unassigned Attributes Assigned Gender Assigned Race
alchemist actress ancient druid
aristocrat ballerino baroque composer
ballet dancer catholic priest colonial governor of cameroon
ceo crusader knight conquistador
craftsman emperor cossack person
detective empress desert nomad
farmer firewoman impi fighter
flight attendant flapper maasai person
housekeeper geisha maharani
human gladiatrix medieval european farmer
interviewer liftboy medieval european peasant
jazz musician maharaja mongolian person
modern philosopher maid olympic medal winner
pirate nun roman legionary
poor person pope samurai
professor president of the united states sheikha
rich person sheikh squire
software developer shieldmaiden suffragette
soldier wehrmacht soldier tsar
sumo wrestler witch viking